OpenStudy (anonymous):

What is the solution to the equation ? -3(h+5)+2=4(h+6)-9

OpenStudy (anonymous):

Solution? Do you mean what does the variable stand for or whether its equal or not?

OpenStudy (anonymous):

-3h -15 +2 = 4h + 24 - 9 -3h -13 = 4h + 15 7h = 28 h =4

OpenStudy (anonymous):

yea what h equals. thanks again Rohith15